OpenStudy (anonymous):

Find the slope of the line by solving for y -5x + 7y = 2 Please show all steps

OpenStudy (saifoo.khan):

-5x + 7y = 2 \[7y = 2 + 5x\] \[y = \frac27+\frac57x\] \[y = \frac37x + \frac27\]

OpenStudy (saifoo.khan):

Slope = 3/7

OpenStudy (anonymous):

THANKS SOOO MUCH

OpenStudy (saifoo.khan):

No problem! :)

OpenStudy (saifoo.khan):

please hit good answer.

OpenStudy (saifoo.khan):

Thanks! :)

OpenStudy (anonymous):

wait how did you get the steps??

OpenStudy (saifoo.khan):

Which part u dont understand?

OpenStudy (anonymous):

how did you get to \[7y = 2 + 5x\]

OpenStudy (saifoo.khan):

added +5x on both sides.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

could you please help with the question i posted before this?